OpenStudy (anonymous):

The polynomial of degree 4, P(x) has a root of multiplicity 2 at x=2 and roots of multiplicity 1 at x=0 and x=−3. Its lead coefficient is 4. Find a formula for P(x).

OpenStudy (mertsj):

Would you agree that if x = 2 then x-2 is a factor?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So, if the root 2 has multiplicity 2, then x-2 must be a factor twice.

OpenStudy (anonymous):

so (x-2)(x-2)(x+4)?

OpenStudy (mertsj):

And if x = 0 is a root, then wouldn't x-0 be a factor?

OpenStudy (mertsj):

And if x = -3 is a root wouldn't x+3 be a factor?

OpenStudy (anonymous):

yea you are right

OpenStudy (anonymous):

so what do i do then? FOIL them?

OpenStudy (mertsj):

So write the factors that you know so far.

OpenStudy (anonymous):

(x-2)(x-2)(x+3)(x-0)

OpenStudy (mertsj):

And that product would be 0, do you agree?

OpenStudy (anonymous):

im not sure. how so?

OpenStudy (mertsj):

Solve this equation: (x-6)(x+9)=0

OpenStudy (anonymous):

x=-9 x=6

OpenStudy (mertsj):

So working backwards, we see that if x = -9 is an answer then x+9 is a factor. Similarly, if x = 6 is an answer then x-6 is a factor and the product has to be 0 based on the 0 factor property

OpenStudy (mertsj):

So your equation must be 4(x-2)(-2)(x)(x+3)=0

OpenStudy (mertsj):

Whether or not you multiply it out depends on what your instructor told you.